### v2

@@ 3.3 Stein Variational Gradient Descent (SVGD)
$\phi^{*}_{k}\propto\mathbb{E}_{x\sim q}[\mathcal{P}k(x,\cdot)]=\mathbb{E}_{x\sim q}\left[\nabla_{x}\log p(x)k(x,\cdot)+\nabla_{x}k(x,\cdot)\right]$ (6)
4 Pareto Set Learning via Stein Variational Gradient Descent
4.1 Stein Variational Hypernetwork
@@
In this study, we develop a novel method for Pareto set learning, which employs Stein Variational Gradient Descent to approximate the entire Pareto solution in Multi-objective learning problems. Furthermore, we explore the different gradient descents for learning in SVGD, providing a complete sense of learning the Pareto set. We name our method SVH-MOL.
We use hypernetwork to generate a set of initial particles, which is an advantage to flexibly changing the number of particles. Within the SVGD framework, each particle’s movement is influenced by the repulsive forces from its neighboring particles. As the particle count increases, these repulsive interactions intensify, leading to faster convergence and a more diverse distribution of solutions across the Pareto front. Additionally, in multi-task learning, rather than training a separate network for every particle, the hypernetwork generates all necessary models on demand. This strategy not only streamlines the overall training process but also significantly reduces memory usage by enabling parameter sharing across tasks.
Building upon the previous study [36], SVH-MOL needs to sample a set of an initial particle by using a hypernetwork. For the input of the hypernetwork, we use the predefined preference vector as a trade-off weight between objectives. Firstly, we sample $n$ preference vectors $r_{i}$ and fit them to the hypernetwork $\mathbf{h}(\cdot,\phi)$ to generate solutions $\{\mathbf{x}_{r_{i}}\}_{i=1}^{n}$. For a problem with $m$ objectives, the objective functions are represented as $\mathcal{F}_{i}(\mathbf{x}_{r_{i}})=[f_{1}(\mathbf{x}_{r_{i}}),...,f_{n}(\mathbf{x}_{r_{i}})]$, where $\mathcal{F}_{i}$ is considered as a particle. The set of particles $\mathbf{F}={\{\mathcal{F}_{i}}\}^{n}_{i=1}$ serves as the initialization set. We then iteratively advance this set of particles towards the Pareto front by applying the update rule in Formula 3.
Stein Variational Gradient Descent plays a role in adjusting the direction of the gradient update and helps push the points away from each other, thereby enhancing the diversity of the Pareto front. Consequently, the following modifications will be applied to Formula 3:
$\phi_{t+1}=\phi_{t}-\xi\sum_{i=1}^{n}\sum_{j=1}^{n}\nabla_{\phi}\mathbf{g}\left(\mathcal{F}_{i}\right)\mathbf{k}(\mathcal{F}_{i},\mathcal{F}_{j})-\alpha\nabla_{\phi}\mathbf{k}(\mathcal{F}_{i},\mathcal{F}_{j})$ (7)
@@
where $\mathbf{k}(\mathcal{F}_{i},\mathcal{F}_{j})$ is kernel matrix. Gaussian kernel is used: $\mathbf{k}(\mathcal{F}_{i},\mathcal{F}_{j})=\exp(-\frac{1}{2h^{2}}||\mathcal{F}_{i}-\mathcal{F}_{j}||^{2})$, $h$ is bandwidth.
  • When considering $\nabla_{\phi}\mathbf{g}\left(\mathcal{F}_{i}\right)\mathbf{k}(\mathcal{F}_{i},\mathcal{F}_{j})$, it is the driving force term that ensures the gradient update direction aligns with the intended path toward the optimal region.
@@
  • The second term, a repulsive force $\nabla_{\phi}\mathbf{k}(\mathcal{F}_{i},\mathcal{F}_{j})$ is responsible for maintaining diversity in the set of particles under consideration.
4.2 Controllable Pareto Set Learning
@@
In traditional Stein Variational Gradient Descent (SVGD), a discrepancy measure known as the Stein discrepancy is used to quantify the difference between the current and target distributions. In a more recent study, [31] introduced Mean Discrepancy Descent (MMD descent), a novel approach that replaces the original driving force in SVGD with one based on the Maximum Mean Discrepancy. Their work highlights a key limitation of SVGD, variance collapse, when approximating high-dimensional distributions. To address this issue, they modify the gradient update formulation, suggesting that for different problems, it may be beneficial to design a problem-specific driving force to guide particles toward better optimal solutions.
@@
In the context of multi-objective learning, the Stein Variational Hypernetwork (SVH) is proposed to approximate the entire Pareto set. This approach leverages Stein’s updating rule while utilizing a hypernetwork to generate candidate solutions. The idea of using a learnable hypernetwork to produce the Pareto set was introduced by [29], who proposed the Controllable Pareto Front Learning (CPF) framework. CPF trains a single hypernetwork whose parameters are optimized using principles from scalarization theory.
@@
To overcome the non-smooth nature of the standard Tchebyshev function, the Smooth Tchebyshev function [14] replaces the max operator with a smooth approximation, such as the softmax or log-sum-exp function. A commonly used form is:
$g(\mathcal{F}_{i})=\frac{1}{\mu}\log\left(\sum_{j=1}^{m}\exp\left(\mu r_{i,j}|f_{j}(x_{r_{i}})-z_{j}^{*}|\right)\right)$ (10)
@@
4.3 Diversity-Promoting via Annealed SVH-MOL
In our optimization process, the gradient updates are influenced by two conflicting components. The first term acts as a driving force, guiding the solutions toward the optimal region. In contrast, the second term serves as a repulsive force, promoting diversity by preventing the solutions from collapsing onto one another along the Pareto front. This issue highlights the trade-off between convergence and diversity. While a Pareto front may exhibit good convergence, it does not necessarily guarantee diversity. In our experiments, we observed that prioritizing the repulsive term often leads to a Pareto set with poor optimality (Figure 2). In previous work, [35] discussed the hyperparameter $\alpha$ in Formula (7), highlighting its influence on the diversity of the Pareto front. However, they were unable to effectively control this effect. In a recent study, [32] also investigated this challenge and reported a variance collapse issue. They observed that particles in SVGD tend to collapse into a few local modes, a phenomenon strongly influenced by the initial particle distribution. To address this, they proposed an annealing schedule that gradually increases the weight of the driving force term. This strategy encourages the initial particles to spread across diverse regions at the beginning, guiding them toward a more diverse distribution throughout the optimization process.
$\phi_{t+1}=\phi_{t}-\xi\sum_{i=1}^{n}\sum_{j=1}^{n}\gamma(t)\nabla_{\phi}\mathbf{g}\left(\mathcal{F}_{i}\right)\mathbf{k}(\mathcal{F}_{i},\mathcal{F}_{j})-\alpha\nabla_{\phi}\mathbf{k}(\mathcal{F}_{i},\mathcal{F}_{j})$ (11)
Where $\gamma(t)$ is a coefficient that controls the influence of the driving term on the overall gradient update. [32] varies $\gamma(t)$ in the interval $[0,1]$ with an appropriate schedule; the process consists of two phases. The first is an exploratory phase, where a predominant repulsive force drives the points away from their initial positions, promoting broad coverage of the target distribution’s support. The second is an exploitative phase, in which the driving force becomes dominant, guiding the points to concentrate around the various modes of the distribution.
Figure 2: Illustration of the change of Pareto front when changing $\alpha$ value.
@@ 4.3 Diversity-Promoting via Annealed SVH-MOL
5 Experiments
5.1 Evaluation Metrics
Mean Euclidean Distance (MED).[29] To evaluate the effectiveness of the controllable Pareto front, which means that the hypernetwork generates the optimal Pareto solution, aligning with the preference vector. We measure the error between the obtained Pareto solutions by the hypernetwork and the true Pareto solutions through calculating the mean of the Euclidean distance. The metric MED is defined as:
@@ 5.1 Evaluation Metrics
Hypervolume (HV). Hypervolume [37] is a metric to figure out both convergence and diversity in the obtained solution. For some problems for which we lack the ground truth Pareto set, HV is an effective way to validate the quality of obtained solutions. The higher HV value reflects that the method is better. Given a set of $n$ points $y=\{y^{(i)}|y^{(i)}\in\mathbb{R}^{m};i=1,\dots,n\}$ and a reference point $\rho\in\mathbb{R}^{m}$, the Hypervolume of $y$ is measured by the region of non-dominated points bounded above by $y^{(i)}\in y$, then the hypervolume metric is defined as follows:
$HV(y)=VOL\left(\underset{y^{(i)}\in y,y^{(i)}\prec\rho}{\bigcup}\displaystyle{\Pi_{i=1}^{n}}\left[y^{(i)},\rho_{i}\right]\right)$ (14)

### v3

@@ 3.3. Stein Variational Gradient Descent (SVGD)
(6) $\phi^{*}_{k}\propto\mathbb{E}_{x\sim q}[\mathcal{P}k(x,\cdot)]=\mathbb{E}_{x\sim q}\left[\nabla_{x}\log p(x)k(x,\cdot)+\nabla_{x}k(x,\cdot)\right]$
4. Pareto Set Learning via Stein Variational Gradient Descent
4.1. Stein Variational Hypernetwork
@@
In this study, we develop a method for Pareto set learning that employs Stein Variational Gradient Descent to approximate the entire Pareto solution in multi-objective learning problems. Our method, SVH-MOL, bridges the theoretical framework of SVGD with practical multi-objective optimization through hypernetworks.
Theoretical Foundation and Connection to SVGD
In traditional SVGD, particles are updated to approximate a target distribution $p(x)$ by minimizing the KL divergence. In our multi-objective context, we reinterpret this framework by considering the Pareto front as the target distribution in objective space. While the Pareto front is not a probability distribution in the conventional sense, we can define a target energy function that captures the desired properties of Pareto optimal solutions.
For a given preference vector $r$, we define the target energy function as:
(7) $E(\mathcal{F})=\mathbf{g}(\mathcal{F},r)$
where $\mathbf{g}$ is the scalarization function. The corresponding unnormalized target density becomes:
(8) $p(\mathcal{F})\propto\exp(-E(\mathcal{F}))=\exp(-\mathbf{g}(\mathcal{F},r))$
This formulation allows us to apply SVGD principles to push particles toward regions of low energy (good Pareto solutions) while maintaining diversity through repulsive forces.
Hypernetwork-based Particle Generation
We employ a hypernetwork to generate initial particles, providing flexibility in adjusting particle count. Within the SVGD framework, each particle’s movement is influenced by repulsive forces from neighboring particles. As particle count increases, these repulsive interactions intensify, leading to faster convergence and more diverse solution distribution across the Pareto front. Building upon (Nguyen_Dinh_Nguyen_Hoang_Le_2025), SVH-MOL samples initial particles using a hypernetwork with predefined preference vectors as trade-off weights between objectives. We sample $n$ preference vectors $r_{i}$ and feed them to the hypernetwork $\mathbf{h}(\cdot,\phi)$ to generate solutions $\{\mathbf{x}_{r_{i}}\}_{i=1}^{n}$. For problems with $m$ objectives, the objective functions are represented as $\mathcal{F}_{i}(\mathbf{x}_{r_{i}})=[f_{1}(\mathbf{x}_{r_{i}}),...,f_{m}(\mathbf{x}_{r_{i}})]$, where each $\mathcal{F}_{i}$ is considered a particle.
The SVGD update rule modifies the hypernetwork parameter update as follows:
(9) $\phi_{t+1}=\phi_{t}-\xi\sum_{i=1}^{n}\sum_{j=1}^{n}\nabla_{\phi}\mathbf{g}\left(\mathcal{F}_{i},r_{i}\right)\mathbf{k}(\mathcal{F}_{i},\mathcal{F}_{j})-\alpha\nabla_{\phi}\mathbf{k}(\mathcal{F}_{i},\mathcal{F}_{j})$
@@
where $\nabla_{\phi}\mathbf{g}(\mathcal{F}_{i},r_{i})=\frac{\partial\mathbf{g}(\mathcal{F}_{i},r_{i})}{\partial\mathcal{F}_{i}}\cdot\frac{\partial\mathcal{F}_{i}}{\partial\phi}$ represents the chain rule through the hypernetwork, and $\mathbf{k}(\mathcal{F}_{i},\mathcal{F}_{j})$ is the Gaussian kernel matrix.
  • The term $\nabla_{\phi}\mathbf{g}(\mathcal{F}_{i})\mathbf{k}(\mathcal{F}_{i},\mathcal{F}_{j})$ acts as the driving force, guiding solutions toward optimal regions by following the gradient of the energy function.
@@
  • The term $\nabla_{\phi}\mathbf{k}(\mathcal{F}_{i},\mathcal{F}_{j})$ serves as the repulsive force, maintaining diversity by preventing particle collapse.
This formulation ensures that solutions simultaneously converge to the Pareto front while maintaining good distribution across it.
4.2. Controllable Pareto Set Learning
@@
In traditional Stein Variational Gradient Descent (SVGD), a discrepancy measure known as the Stein discrepancy is used to quantify the difference between the current and target distributions. In a more recent study, (ba2021understanding) introduced Mean Discrepancy Descent (MMD descent), a novel approach that replaces the original driving force in SVGD with one based on the Maximum Mean Discrepancy. Their work highlights a key limitation of SVGD, variance collapse, when approximating high-dimensional distributions. To address this issue, they modify the gradient update formulation, suggesting that for different problems, it may be beneficial to design a problem-specific driving force to guide particles toward better optimal solutions.
@@
In the context of multi-objective learning, the Stein Variational Hypernetwork (SVH) is proposed to approximate the entire Pareto set. This approach leverages Stein’s updating rule while utilizing a hypernetwork to generate candidate solutions. The idea of using a learnable hypernetwork to produce the Pareto set was introduced by (tuan2024framework), who proposed the Controllable Pareto Front Learning (CPF) framework. CPF trains a single hypernetwork whose parameters are optimized using principles from scalarization theory.
@@
To overcome the non-smooth nature of the standard Tchebyshev function, the Smooth Tchebyshev function (lin2024smooth) replaces the max operator with a smooth approximation, such as the softmax or log-sum-exp function. A commonly used form is:
(12) $\mathbf{g}(\mathcal{F}_{i},r_{i})=\frac{1}{\mu}\log\left(\sum_{j=1}^{m}\exp\left(\mu r_{i,j}|f_{j}(x_{r_{i}})-z_{j}^{*}|\right)\right)$
@@
4.3. Diversity-Promoting via Annealed SVH-MOL
Figure 2. Illustration of the trade-off controlled by the diversity hyperparameter $\alpha$ (see Eq. 11). A large $\alpha$ (red dots) prioritizes the repulsive force, leading to a diverse but suboptimal Pareto set (poor convergence). A small $\alpha$ (black dots) results in a well-converged front but with potentially less spread.
In our optimization process, the gradient updates are influenced by two conflicting components. The first term acts as a driving force, guiding the solutions toward the optimal region. In contrast, the second term serves as a repulsive force, promoting diversity by preventing the solutions from collapsing onto one another along the Pareto front. This issue highlights the trade-off between convergence and diversity. While a Pareto front may exhibit good convergence, it does not necessarily guarantee diversity. In our experiments, we observed that prioritizing the repulsive term often leads to a Pareto set with poor optimality (Figure 2). In previous work, (liu2021profiling) discussed the hyperparameter $\alpha$ in Formula (9), highlighting its influence on the diversity of the Pareto front. However, they were unable to effectively control this effect. In a recent study, (d2021annealed) also investigated this challenge and reported a variance collapse issue. They observed that particles in SVGD tend to collapse into a few local modes, a phenomenon strongly influenced by the initial particle distribution. To address this, they proposed an annealing schedule that gradually increases the weight of the driving force term. This strategy encourages the initial particles to spread across diverse regions at the beginning, guiding them toward a more diverse distribution throughout the optimization process.
(13) $\phi_{t+1}=\phi_{t}-\xi\sum_{i=1}^{n}\sum_{j=1}^{n}\gamma(t)\nabla_{\phi}\mathbf{g}\left(\mathcal{F}_{i}\right)\mathbf{k}(\mathcal{F}_{i},\mathcal{F}_{j})-\alpha\nabla_{\phi}\mathbf{k}(\mathcal{F}_{i},\mathcal{F}_{j})$
Where $\gamma(t)$ is a coefficient that controls the influence of the driving term on the overall gradient update. (d2021annealed) varies $\gamma(t)$ in the interval $[0,1]$ with an appropriate schedule; the process consists of two phases. The first is an exploratory phase, where a predominant repulsive force drives the points away from their initial positions, promoting broad coverage of the target distribution’s support. The second is an exploitative phase, in which the driving force becomes dominant, guiding the points to concentrate around the various modes of the distribution.
@@ 4.3. Diversity-Promoting via Annealed SVH-MOL
Algorithm 1 Annealed Stein Variational Hypernetwork for Multi-Objective Learning (A-SVH-MOL)
1:Objective functions $\mathcal{F}=\{f_{1},\ldots,f_{m}\}$; number of particles $n$; learning rate $\xi$; diversity hyperparameter $\alpha$; annealing parameters $T_{0}$, $\tau$; number of training epochs $E$
2:Trained hypernetwork parameters $\phi^{*}$
3:Initialize hypernetwork parameters $\phi$
4:for epoch $t=1$ to $E$ do
5:  Sample batch of $n$ preference vectors $\{r_{i}\}_{i=1}^{n}\sim\text{Dir}(\alpha)$
6:  Generate particles via hypernetwork: $\{x_{i}\}_{i=1}^{n}=h(\{r_{i}\}_{i=1}^{n},\phi_{t})$
7:  Compute objective values: $\{F_{i}\}_{i=1}^{n}=\{\mathcal{F}(x_{i})\}_{i=1}^{n}$
8:  Calculate annealing coefficient $\gamma(t)$ using Eq. (14)
9:  Initialize gradient accumulator: $\nabla_{\phi}I\leftarrow 0$
10:  for each particle $i=1$ to $n$ do
11:   Compute $\nabla_{\phi}\mathbf{g}(\mathcal{F}_{i})$ $\triangleright$ Gradient of scalarization function w.r.t. $\phi$
12:   for each particle $j=1$ to $n$ do
13:     Compute kernel: $k_{ij}=\exp\left(-\frac{\|F_{i}-F_{j}\|^{2}}{2h^{2}}\right)$
14:     Compute driving force: $\text{drive}=\gamma(t)\cdot\nabla_{\phi}\mathbf{g}(\mathcal{F}_{i})\cdot k_{ij}$
15:     Compute repulsive force: $\text{repulse}=\alpha\cdot\nabla_{\phi}k_{ij}$
16:     Accumulate gradient: $\nabla_{\phi}I\leftarrow\nabla_{\phi}I+\text{drive}+\text{repulse}$
17:   end for
18:  end for
19:  Update parameters: $\phi_{t+1}\leftarrow\phi_{t}-\xi\cdot\nabla_{\phi}I$
20:end for
21:return $\phi^{*}\leftarrow\phi_{E}$
5. Experiments
5.1. Evaluation Metrics
Mean Euclidean Distance (MED).(tuan2024framework) To evaluate the effectiveness of the controllable Pareto front, which means that the hypernetwork generates the optimal Pareto solution, aligning with the preference vector. We measure the error between the obtained Pareto solutions by the hypernetwork and the true Pareto solutions through calculating the mean of the Euclidean distance. The metric MED is defined as:
@@ 5.1. Evaluation Metrics
Hypervolume (HV). Hypervolume (zitzler1999multiobjective) is a metric to figure out both convergence and diversity in the obtained solution. For some problems for which we lack the ground truth Pareto set, HV is an effective way to validate the quality of obtained solutions. The higher HV value reflects that the method is better. Given a set of $n$ points $y=\{y^{(i)}|y^{(i)}\in\mathbb{R}^{m};i=1,\dots,n\}$ and a reference point $\rho\in\mathbb{R}^{m}$, the Hypervolume of $y$ is measured by the region of non-dominated points bounded above by $y^{(i)}\in y$, then the hypervolume metric is defined as follows:
(16) $HV(y)=VOL\left(\underset{y^{(i)}\in y,y^{(i)}\prec\rho}{\bigcup}\displaystyle{\Pi_{i=1}^{n}}\left[y^{(i)},\rho_{i}\right]\right)$